Fermi National Laboratory


Accelerator Update - June 18-June 20

Wednesday June 18
The day shift began with the Antiproton Source (Pbar) stacking, with Tevatron (TeV) experts conducting beam maintenance studies, with Recycler experts conducting studies, with Switchyard studies, with Booster experts tuning the Booster beam, and with MiniBooNE taking beam.

A power supply expert reported that a Pbar Debuncher SEPTUM (D:ISEPV) was making noise. He was turning the device off to investigate at 8:53 AM. He repaired a bad cable and had the power supply back on by 9:55 AM. Pbar resumed stacking.

At 10:45 AM, a Linac RF station (KRF4) tripped off and would not reset. A Linac technician had the station operating ten minutes later. Then at 11:44 AM, a Linac quadrupole power supply started causing problems. An expert investigated and repaired this also.

Switchyard experts sent beam to the Switchyard dump while the TeV was at 980 GeV. The TeV then ramped back down to 150 GeV without problems. (This was an important step in proving that Switchyard and the TeV could work simultaneously without affecting each other.)

TeV experts completed their maintenance studies at 12:30 PM. Operators prepared the Collider Detector Facility (CDF) and D-zero (D0) collision halls for access. Both experiments were out of access by 2:30 PM.

Switchyard experts completed their studies at 3:22 PM.

Operations put the TeV through a dry squeeze at 3:39 PM.

Operation put the TeV into shot setup at 4:52 PM.

Operations established store 2701 at 8:14 PM with an initial luminosity of 28.55E30. Pbar resumed stacking.

A NuMI expert began a study period at 8:45 PM.

At 11:03 PM, all but 5 Main Injector (MI) RF stations tripped off. MiniBooNE also tripped off. These trips were due to NuMI studies. Everything reset okay. But then it happened again at 11:23 PM. Again everything reset okay.

Operators took a MI RF station (MIRF14) that had failed out of the system.

Thursday June 19
The midnight shift began with Operations monitoring store 2701, with Pbar stacking, and with MiniBooNE taking beam.

At 1:27 AM, the TeV suffered a sector A quench on the dipole corrector magnets located there. The cause may have been due to a quench protection card failure. This card was later replaced.

Booster suffered a vacuum burst on RF stations 11 and 12.

A Pbar Accumulator kicker (A:IKIK) tripped off at 3:41 AM. Operations called in an expert. He had the kicker back on by 4:47 AM.

MI beam aborted due to a MECAR problem.

At 7 AM, the NuMI project suffered a power outage. This appeared to be a problem on a line that feeds only NuMI. A high voltage expert investigated and discovered the remains of a raccoon that had caused fuses to blow.

Preaccelerator experts began work on the H- Source at 11:21 AM. The work was completed fifteen minutes later.

Operations put the TeV into shot set up and at 11:53 AM established store 2705 with an initial luminosity of 22E30. Pbar resumed stacking.

Vacuum techs replaced vacuum pumps for BRF13 and 18 at 3:50 PM.

Booster extraction tuning was still continuing at 8:12 PM.

A MI expert began multi-batch studies at 9:30 PM.

At 10:48 PM, NuMI studies caused more MI trips. Operators had things back to normal within thirty minutes.

Friday June 20
The midnight shift began with Operations monitoring store 2702, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 2702 at 1 AM.

At 1:23 AM, the CDF Roman Pots decided, on their own, to take a stroll into the colliding beam. All but two of them were stopped and driven back out. On issuing a second command to drive out, these two pots moved out, but another decided to move back in. Operators began changing out control cards. Eventually, all the pots were driven out and then disabled. Experts will work on this later in the day.

Operations put the TeV into a dry squeeze at 1:30 AM.

Operations put the TeV into shot setup at 1:58 AM.

At 3:13 AM, something odd happened with the antiproton transfers for the next store; immediately after transfer 5 went into the TeV another transfer occurred and was lost. There was no order for this transfer. Operators called in an expert.

Operations established store 2707 at 3:38 PM with an initial luminosity of 27.8E30.

While trying to resume stacking, at 3:55 AM, a control card failed. Operators had the card replaced by 4:39 AM.

Though problems existed with Pbar Debuncher (DRF1) phase loops, stacking resumed at 3:57 AM.

MI had a vacuum burst at 7 AM, and then at 7:37 AM two RF stations (BRF5 and 6) tripped off due to vacuum problems. An expert was called to investigate.

Today's Plan
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/23/2003   email Fermilab